OpenStudy (anonymous):

Which point is the vertex of the parabola y = x2+4x+1?

OpenStudy (anonymous):

y = x² + 4x + 1 Vertex ( h, k) h= -b/2a, = -4/ 2 = -2 y = f( -b/2a) f(-2) = 4 -8 + 1 = -3 => Vertex ( -2, -3)

OpenStudy (anonymous):

thank you (:

OpenStudy (anonymous):

Any questions?

OpenStudy (anonymous):

as of now no i just need 2 more questions to be answered

OpenStudy (anonymous):

same topic

OpenStudy (anonymous):

Just keep posting more and more questions, you'll learn in no time :)

OpenStudy (anonymous):

well thank you (: